OpenStudy (anonymous):

Identify the vertical asymptotes of f(x) = quantity x minus 4 over quantity x squared plus 13 x plus 36

OpenStudy (anonymous):

\[x - 4 \over x^2 + 13x + 36\]

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

well you need to so a little work in factoring as the vertical asymptotes are the values of x, that make the denominator zero. Remember you can't divide by zero, its undefined. so factoring you have (x +4)(x + 9) = 0 or x + 4 = 0 x + 9 = 0 solve both of these equations and you'll have the vertical asymptotes...

OpenStudy (anonymous):

brb for one sec...

OpenStudy (anonymous):

x = -4 and x = -9 lol?

OpenStudy (campbell_st):

yep... thats it...

OpenStudy (anonymous):

sweet